Q1: Use the matrix

9 1

7 2 

(a)  Obtain the Hill cipher for the plaintext message

                                              PAK ARMY

by letting  A=1,B=2,C=3,.......,Y=25 and Z=26

(b)  Decode Hill 2- cipher which was encrypted by this matrix.


1
Expert's answer
2021-11-29T08:50:41-0500

To encrypt a message, each block of n letters (considered as an n-component vector) is multiplied by an invertible n × n matrix, against modulus 26. To decrypt the message, each block is multiplied by the inverse of the matrix used for encryption.


a)

"PAK\ ARMY\to \begin{pmatrix}
 P \\
 A
\end{pmatrix}, \begin{pmatrix}
 K \\
 A
\end{pmatrix}, \begin{pmatrix}
 R \\
 M
\end{pmatrix}, \begin{pmatrix}
 Y \\
 X
\end{pmatrix}\to\begin{pmatrix}
 16 \\
 1
\end{pmatrix}, \begin{pmatrix}
 11 \\
 1
\end{pmatrix}, \begin{pmatrix}
 18 \\
 13
\end{pmatrix}, \begin{pmatrix}
 25 \\
 24
\end{pmatrix}"


"\begin{pmatrix}
 9 & 1 \\
 7 & 2
\end{pmatrix}\begin{pmatrix}
 16 \\
 1
\end{pmatrix}=\begin{pmatrix}
 145 \\
 114 
\end{pmatrix}=\begin{pmatrix}
 15 \\
 10 
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 9 & 1 \\
 7 & 2
\end{pmatrix}\begin{pmatrix}
 11 \\
 1
\end{pmatrix}=\begin{pmatrix}
 100 \\
 79 
\end{pmatrix}=\begin{pmatrix}
 22 \\
 1 
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 9 & 1 \\
 7 & 2
\end{pmatrix}\begin{pmatrix}
 18 \\
 13
\end{pmatrix}=\begin{pmatrix}
 175 \\
 152 
\end{pmatrix}=\begin{pmatrix}
 19 \\
 22 
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 9 & 1 \\
 7 & 2
\end{pmatrix}\begin{pmatrix}
 25 \\
 24
\end{pmatrix}=\begin{pmatrix}
 249 \\
 199 
\end{pmatrix}=\begin{pmatrix}
 15 \\
 17
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 15 \\
 10
\end{pmatrix}, \begin{pmatrix}
 22 \\
 1
\end{pmatrix}, \begin{pmatrix}
 19 \\
 22
\end{pmatrix}, \begin{pmatrix}
 15 \\
 17
\end{pmatrix}\to \begin{pmatrix}
 O \\
 J
\end{pmatrix}, \begin{pmatrix}
 V \\
 A
\end{pmatrix}, \begin{pmatrix}
 R \\
 M
\end{pmatrix}, \begin{pmatrix}
 X \\
 Y
\end{pmatrix}"


b)

"\begin{pmatrix}
 9 & 1 \\
 7 & 2
\end{pmatrix}^{-1}=\frac{1}{11}\begin{pmatrix}
 2 & -1 \\
 -7 & 9
\end{pmatrix}\equiv 19\begin{pmatrix}
 2 & -1 \\
 -7 & 9
\end{pmatrix}=\begin{pmatrix}
 12 & 25 \\
 19 & 15
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 12 & 25 \\
 19 & 15
\end{pmatrix}\begin{pmatrix}
 15 \\
 10
\end{pmatrix}=\begin{pmatrix}
 430 \\
 435
\end{pmatrix}=\begin{pmatrix}
 16 \\
 1
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 12 & 25 \\
 19 & 15
\end{pmatrix}\begin{pmatrix}
 22 \\
 1
\end{pmatrix}=\begin{pmatrix}
 279 \\
 433
\end{pmatrix}=\begin{pmatrix}
 11 \\
 1
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 12 & 25 \\
 19 & 15
\end{pmatrix}\begin{pmatrix}
 19 \\
 22
\end{pmatrix}=\begin{pmatrix}
 778 \\
 691
\end{pmatrix}=\begin{pmatrix}
 18 \\
 13
\end{pmatrix}(mod\ 26)"


"\begin{pmatrix}
 12 & 25 \\
 19 & 15
\end{pmatrix}\begin{pmatrix}
 15 \\
 17
\end{pmatrix}=\begin{pmatrix}
 605 \\
 540
\end{pmatrix}=\begin{pmatrix}
 25 \\
 24
\end{pmatrix}(mod\ 26)"